OpenStudy (xxharleyquinnxx):

the distance from the tetherball pole to the edge of the circle is 5 feet what is the circumference of the circle

OpenStudy (tshimp0629):

ummm you do know this is the history section right?

OpenStudy (hope210):

Please post this in the math section :) --> http://openstudy.com/study#/groups/Mathematics

OpenStudy (maddisongruby):

The equasion your looking for is C=pi* D you are given the radius, diameter= r*2 so in this case D=5*2 so D=10 now plug in D with 10 C=10* pi that gives you roughly C=31.42ft